In an old house, the heating system uses radiators, which are hollow metal devices through which hot water or steam circulates. In one room the radiator has a dark color (emissitivity = 0.820). It has a temperature of 60.4 oC. The new owner of the house paints the radiator a lighter color (emissitivity = 0.491). Assuming that it emits the same radiant power as it did before being painted, what is the temperature (in degrees Celsius) of the newly painted radiator?
1
Expert's answer
2020-09-22T17:10:04-0400

Radiant power, emitted by the body with area "A" , temperature "T" and emissivity "\epsilon" is

"P = Aj= \epsilon \sigma T^4 A" , where "j= \epsilon \sigma T^4" is Stefan-Bolzmann law.

According to the task, "P_1 = P_2"

"\epsilon_1 \sigma T_1^4 A = \epsilon_2 \sigma T_2^4 A"

"\epsilon_1 T_1^4 = \epsilon_2 T_2^4"

Converting from Celsius to Kelvin,

"T_1 = t _1+ 273.15 = 60.4 + 273.15 = 333.55\, K"

"\displaystyle T_2 = (\frac{\epsilon_1}{\epsilon_2})^{\frac{1}{4}} T_1= (\frac{0.820}{0.491})^{0.25} 333.55 =1.1368 \cdot 333.55 =379.18 \; K"

"t_2 = T_1 - 273.15 = 106.03^\circ C"

Answer: "t_2 = 106.03^\circ C"